OpenStudy (hari5719):

CAN U PLS HELP ME WITH THIS MATH QUESTION!!!1

OpenStudy (hari5719):

HERE IS A SCREENSHOT

OpenStudy (hari5719):

@ParthKohli @nincompoop @triciaal @ikram002p @goldengoodgirl @jad3

OpenStudy (hari5719):

@mathmale @LegendarySadist

OpenStudy (mathmale):

What have you been able to do so far? What have you learned very recently that relates to this type of question?

OpenStudy (mathmale):

Hari: I need your involvement in order to help you. I'm moving on. You might find "Pascal's Triangle" to be useful here.

OpenStudy (hari5719):

yes its binomial theorem i suppose @mathmale

imqwerty (imqwerty):

if you got something like this-\((a+b)^n\) just write the general term and it is given like this- \(\Large T_{n+1} = ^nC_r (a)^{n-r}(b)^r \) 1st write the general term for your expression and then the equate the power of x to the desired power u'll get r and then when you get r put it in your general term equation and simplify it then you can get the coefficient.

imqwerty (imqwerty):

this^ is for part 1 for part 2 you gotta write the general term for the \(\large \left( 2x +\frac{1}{2x} \right)^6\) part and then just multiply the general term with \((1+x^2)\) then you have to evaluate and find the values of r for which you get ther \(x^2\) term find all \(r\) and then find the terms and add them and get the coefficient

OpenStudy (mathmale):

Pascal's Triangle gives us a fast way to carry out binomial expansions, by which I mean multiplying out expressions such as (a+b)^3 = (a+b)(a+b)(a+b). Pascal's Triangle generates the constant coefficients of such expansions. Suppose we want to do an actual example: Find the binomal expansion of (a+b)^3. Write out the first four rows of Pascal's Triangle, as follows: